Journalist Report – Nov 10th

Sun 10 Nov Sol 0
by Guy Murphy

Today marked the completion of the changeover between Crews 214 and 215, with Andrew Wheeler taking over as Crew Commander. The first fortnight of Expedition Boomerang has been different from most crew rotations, in that it comprised 4 people. Early Mars mission plans generally call for crew sizes of between 4-8, so this was not an unrealistic scenario, though MDRS crews usually comprise at least 6. In practice, this meant sharing Hab duties and report writing requirements amongst fewer hands, while also trying to complete our research projects over the fortnight.

The crew bonded and worked together extremely well. We rose to a series of unforeseen challenges along the way, including a shortage of food in the first week. Crew 214 demonstrated that a smaller crew can run a mission successfully if it contains the right mix of skills, motivation, and temperaments.

At around 10 am Dianne and Sandy carried their luggage downstairs and farewelled the MDRS after 15 extraordinary days here. Andrew and I are staying on for Crew 215, so we’re very sad to see them leave, but knew they would be soon enjoying the comforts of Earth life back at Grand Junction. Their departure left the 6 team members of Crew 215 together at the campus.

After lunch, Atila and David came over from the outpost to train the new crew members in the use of radios, space suits, and rovers. Shannon Rupert provided further briefings about other issues.

Relative to the beginning of Crew 214, the Hab pantries are now abundantly stocked with a variety of foods, including some treats the new crew members brought up from Grand Junction. Larissa and Jennifer prepared a magnificent spaghetti bolognese for dinner.

Unfortunately, communications were down in the evening (possibly due to another solar storm?), but we were able to submit mission reports via a secondary channel.

Tomorrow morning we will enter full simulation mode. We will wake up on Mars.

Greenhab Report – Nov 10th

Crew 215 GreenHab Report 10-11-2019
GreenHab Officer: Guy Murphy

Environmental control: Door left open for ventilation early the afternoon then reclosed 5 pm.

Average temperatures: –
Hours of supplemental light: Automatic system

Daily water usage for crops: 16 litres (2 full blue watering cans).
Daily water usage for research and/or other purposes:
Water in Blue Tank _____ gallons
Time(s) of watering for crops: 3:35pm.

Changes to crops: Seeds planted described further below.
Established plants introduced to Greenhab 03/11/2019
[3x] strawberry plants (Everbearing), in medium metal planter
[2x] small Aloe Vera, in medium metal planter, frost damaged
[1x] spicy orange thyme (thymus fragantissimus), small terra cotta pot
[1x] lemon balm scented geranium (pelargonium x Melissinum), small terra cotta pot
[1x] Mint Mentha – Berries & Cream, small terra cotta pot
[1x] Pineapple Mint – mentha suaveolens variegated, small terra cotta pot
[1x] Ice plant – Delospermo ‘jewel of the desert garnet’, small terra cotta pot
[1x] Sempervium textorum assorted Hen & Chicks, small terra cotta pot
[1x] Rosemary ‘Barbecue’ (Rosmarinus officinalis ‘Barbecue’), large plastic pot
[1x] Rosemary ‘Spice Island’ Rosmarinus officinalis ‘Spice Island’, large plastic pot
[1x] mint, grapefruit, large plastic pot

Seeds Planted in Greenhab 04/11/2019 (Large metal planting tub)
Greek Oregano – Origanum heracleoticum
Chives – Allium schoenoprasum
Sweet Marjoram – Origanum majorana
Lemon Balm – Melissa officialnis
Sage – Savia Officinalis
Thai Basil – Siam Queen
German/Winter Thyme – Thymus vulgaris
Seeds Planted in Greenhab 10/11/2019 (Large metal planting tub)
Corinader – Coriandrum Sativum
Other seeds stored in the Greenhab but not yet planted are identified as –
Snow Pea (Oregon Giant)
Cucumbers (Patio Snackers F1)
Pepper (Mini Bell Color Mix)
Tomato (Chocolate Cherry)
Perpetual Spinach (Green Leaf Chard)

Harvest: None
Support/supplies needed: N/A

Science Report – Nov 10th

Science Report 10 November 2019
SOL: 0
Crew 215 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler

Science Dome Operations: The Science Dome has been activated and is awaiting EVAs for sample acquisition and testing.

Green Hab: The green hab is fully operational and sown plants, seedlings and established plants are being maintained. See GreenHab report.

Astronomy Dome Operations: No astronomical observations have been planned for this crew rotation.

RAM Operations: The RAM Dome has been activated and is ready for contingencies.

EVA: The first EVA request for geological reconnaissance has been sent. See EVA Request.

EVA Suit Maintenance: The crew engineer will troubleshoot Suit 10 to enable repairs. Suit 2 is experiencing a reduction in airflow and will also be evaluated. See Operations Report.

Additional activities: Nil.

Sol Summary – Nov 10th

Crew 215 Sol Summary Report 10-NOV-2019
Sol: 0
Summary Title: Crew 215 in residence
Author’s name: Andrew Wheeler

Mission Status: In the commissioning phase

Sol Activity Summary: Introducing Crew 215. Andrew Wheeler Crew Commander and Science Officer is an exploration geologist continuing sampling for the micrometeorite flux and use of gypsum as an ISRU analogue. Guy Murphy Green Hab Officer and Journalist is the Mars Society representative who will grow and nurture the crops and provide daily commentary including photos. Steve Whitfield Health and Safety Officer, from Griffith University, is a lecturer in paramedicine who will investigate health and well being in isolated communities. Jennifer Lane EVA coordinator is a Curtin University doctoral candidate in ‘Marsitecture’ and is investigating habitat design. Larissa Wilson Quartermaster of food and water stores is a mechanical engineering student at the University of Canterbury (NZ) evaluating acoustics, will be identifying and attenuating noise sources and levels. Shane Usher Crew Engineer and IT Officer is a chemist and chemical engineer from the University of Melbourne who will be evaluating communications, data management and information dissemination within closed systems. After introductions, later in the morning, we bid farewell to crew 214 members Dianne McGrath and Sandy Dance before spending the afternoon with the hab briefing, doffing and donning of space suits, egress, and ingress through the airlocks and operation of rovers. Finally, after a hot meal, we settled in to await our first dawn on Mars.

Look Ahead Plan: Reconnaissance EVA to the wash area south of Zubrin’s Head. Preparations for the setting out of the micrometeorite sampling grid.

Anomalies in work: N/A

Weather: Marginally below freezing overnight, no breeze, partly cloudy.

Crew Physical Status: Healthy

EVA: None

Reports to be file: Sol Summary Report, Operations Report, HS Report, GreenHab Report, EVA Request, Photos.

Support Requested: Nil

Operations Report – Nov 10th

Crew 214 Operations Report 10-11-2019
SOL: 0
Name of person filing report: Shane Usher

Non-nominal systems: Opportunity, Suit 10, Suit 5 Helmet, No power to multiple downstairs wall sockets.
Notes on non-nominal systems: Opportunity brakes do not work, Suit 10 is not functioning, Suit 5 Helmet falling apart

Generator: run
Hours run: 12h
From what time last night: 1915
To what time this morning: 0729
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night): 82%

Diesel Reading – 55%
Station Propane Reading – 74%
Ethanol Free Gasoline: N/A

Water (loft tank): 23 gallons
Water Meter: 1457085 units
Water (static tank): 150 gallons (not filled)
Static to Loft Pump used – no
Water in Green Hab: 225 gallons
Water in Science Dome: 0 gallons

Toilet tank emptied: yes

Deimos rover used: not at hab
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –

Spirit rover used: yes
Hours: 109.6
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 87%
Currently charging: yes. Note that opening hood to clean mud from engine compartment loosened power cable, that has now been addressed. The power cable was connected to un-powered wall sockets and was an obstruction in hab rear entrance. Cable has been re-routed out of the way to power powerboard above Hab rear air-lock.

Opportunity rover used: no
Hours: 65.6
Beginning charge: 100%
Ending charge: 100%
Currently charging: no

Curiosity rover used: yes
Hours: 115.2
Beginning charge: (Before EVA) : 100%
Ending charge: (On return from EVA, before charging): 100%
Currently charging: yes

Notes on rovers: no
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
ATV Fuel Used: 0 Gals
# Hours the ATVs were used today: 0
Notes on ATVs: Assigned to the director

HabCar used and why, where? Still in Bicknel for service
CrewCar used and why, where?. CrewCar: 9 Nov MDRS Odometer 179069 miles, 9 Nov Refuelled at Hollow Mountain, 9 Nov picked up Crew 215 at Grand Junction 179241, 9 Nov refueled at Hollow Mountain, 9 Nov returned to MDRS, 10 Nov Refuelled at Hollow Mountain, 10 Nov to Grand Junction for Crew 214

General notes and comments: N/A
Summary of internet: Functional
Summary of suits and radios: Suit 2 has poor airflow, Suit 5 Helmet falling apart, Suit 10 is not functioning will investigate and seek guidance from CAPCOM.
Summary of Hab operations: All systems nominal apart from; Stateroom5, Edison screw lightbulb Max 13W required; Fire alarm test button not working in Hab upper level; Fire alarm test button disintegrated in GreenHab.
Summary of GreenHab operations: Atmospheric equalization to the green hab was carried out during the day, plants were watered (16L). The heater automatically came on at 5 pm.
Summary of Science Dome operations: All systems nominal.
Summary of RAM operations: Not used
Summary of any observatory issues: Nominal
Summary of health and safety issues: Shane Usher – mild headache, staying hydrated, 2 panadol

Questions, concerns and requests to Mission Support: Static tank not filled. Water heater unusable. Opportunity is currently unusable.

Crew 215 Journalist Report 19Nov2019

[title Journalist Report – November 19th]

Tues 19 Nov Sol 9
by Guy Murphy
Today was planned against the countdown to rain expected this evening, which will turn
the surrounding landscape of absorbant clays into unpassable slush. The usually
dessicating air is become more humid. Any outdoor activities we hoped to do on this
crew rotation needed to be completed by the end of today. Shane, Larissa and I ventured
out on EVA tracing the entry road towards the edge of the public lands to return some
rock samples to their original context. We stopped to take some final photos of each
other wearing simulated space suits along the way. We have completed our sampling of
magnetic particles undertaken as part of the micro-meteorite study, so this was most
likely the final EVA for Crew 215.
When we were nearest Hanksville, my mobile phone suddenly begun beeping with text
messages. I had momentarily picked up a phone signal for the first time in some weeks.
Whether by design or otherwise, the MDRS campus is located in a mobile phone signal
black spot, and I had not received any text messages while there.
On Mars, communication with home will be strained. The internet protocal has been
extended to communicate with other planets and deep space, and between Earth and Mars
this will involve time delays of between approximately 4 and 24 minutes each way
depending on where the planets are located. Bandwidth will be restricted, so the high
speed, data rich existance people on Earth are accustomed to will not be possible.
Reflecting these constraints, the MDRS has a data limits on its internet connection,
which has to be shared between the crew and also used for submitting formal reports.
This is one of the major lifestyle differences of living here in simulation mode, and
takes some getting used to, but its a chance to rediscover the moment and ponder how
much of our ‘connectedness’ in really necessary. Online video is out of the question,
though new recipes for the bread machine are not.
The lack of phone contact and limited internet is supplemented with direct radio with
the nearby onsite support crew. This evening we are hearing crackling on the radio of
distant approaching storms. A lot of static electricity is generated in this dry
environment, and depending on my clothing, there may be a small crack or flash when I
touch the metal ladder balustrade at night.

Crew 215 Greenhab Report 19Nov2019

[title Greenhab Report – November 19th]

Crew 215 GreenHab Report 19-11-2019
GreenHab Officer: Guy Murphy
Environmental control: Door left open for ventilation midday then reclosed 5pm as the
day was overcast.
Average temperature: 23C (temperature read 12.07pm, 3:12pm)
Hours of supplemental light: Automatic system
Daily water usage for crops: 20 litres (2.5 full blue watering cans).
Daily water usage for research and/or other purposes:
Water in Blue Tank –
Time(s) of watering for crops: 8:30am.
Changes to crops: Planted 25 white Narcissus bulbs in pots.

SEASONAL PLANTING HISTORY
Established plants introduced 03/11/2019

[3x] strawberry plants (Everbearing), in medium metal planter
[1x] spicy orange thyme (thymus fragantissimus), small terra cotta pot
[1x] lemon balm scented geranium (pelargonium x Melissinum), small terra cotta pot
[1x] Mint Mentha – Berries & Cream, small terra cotta pot
[1x] Pineapple Mint – mentha suaveolens variegated, small terra cotta pot
[1x] Ice plant – Delospermo ‘jewel of the desert garnet’, small terra cotta pot
[1x] Sempervium textorum assorted Hen & Chicks, small terra cotta pot
[1x] Rosemary ‘Barbecue’ (Rosmarinus officinalis ‘Barbecue’), large plastic pot
[1x] Rosemary ‘Spice Island’ Rosmarinus officinalis ‘Spice Island’, large plastic pot
[1x] mint, grapefruit, large plastic pot
[2x] small Aloe Vera, in medium metal planter, frost damaged (removed 1311219)

Seeds Planted 04/11/2019 (Large metal planting tub)
Greek Oregano – Origanum heracleoticum
Chives – Allium schoenoprasum
Sweet Marjoram – Origanum majorana
Lemon Balm – Melissa officialnis
Sage – Savia Officinalis
Thai Basil – Siam Queen
German/Winter Thyme – Thymus vulgaris

Seeds Planted 10/11/2019 (Large metal planting tub)
Corinader – Coriandrum Sativum

Seeds Planted in 12/11/2019 (2 smaller planting tubs)
Carrots ‘Purple Sun F1’
Radish ‘icicle’

Seeds Planted 13/11/2019
Perpetual Spinach, Rocket, Spinach ‘Lakeside F1’in medium metal planter.
Tomato (Chocolate Cherry) and Tomato Pole Roma in small pots for transplantation.

Seeds Planted 14/11/2019
Tomato (Chocolate Cherry) and Tomato Pole Roma in small pots for transplantation.
Onion (Evergreen White Bunching)
Cucumbers (Patio Snackers F1) in hanging pot
Pepper (Mini Bell Color Mix) in small pots for transplantation.

Seeds Planted 15/11/2019
Acorn Pumpkin in a small pot for future transplantation. Seed taken from a pumpkin
brought to the Hab at start of rotation as a vegetable.
Snow Peas in 2 hanging pots.

Bulbs planted 19/11/2019
25 white Narcissus bulbs (Paperwhites ‘Ziva’)
Planted into 3 small terra cotta pots & 2 long planter boxes.

Other seeds stored in the Greenhab are identified as –

Rocket ‘Arugula Rocket Salad ‘Rocky’ (Diplotaxis tenuifolia), 1/2 packet
Spinach ‘Lakeside F1’, 1/2 packet
Carrots ‘Purple Sun F1’ – 1/2 packet
Radish ‘icicle’ – 1/2 packet
Snow Pea (Oregon Giant)
Cucumbers (Patio Snackers F1), 1/2 packet
Onion (Evergreen White Bunching), 1/2 packet
Pepper (Mini Bell Color Mix), 1/2 packet
Perpetual Spinach (Green Leaf Chard), 1/2 packet

Partly used packed of seed have been stored in the provided tin to protect from vermin.
This will also also repropogation later in the season to stagger harvest times and
allow resowing if crops are lost.

Harvest: None

Support/supplies needed: Last night 3 bags potting mix, 5 bags garden soil & 2 aloe
vera plants were delivered.

Crew 215 EVA Report 19Nov2019

[title EVA Report – November 19th]

Crew 215 EVA Report 19-Nov-2019
EVA # 10
Author: Jennifer Lane
Purpose of EVA: To return rock samples used for training to collection sites
Start time: 0915
End time: 1025
Narrative: We made our way south on Cow Dung road to road 1101 where we pulled up and
found a nice place to return our agate samples, among other things that we had been
viewing through the microscope. On our way back up Cow Dung Road we stopped to take
photos at several locations, noting that this was likely to be our last EVA due to the
incoming inclement weather.
Destination: Entry to road 1101
Coordinates (use UTM NAD27 CONUS): 519500E 4247300N
Participants: Guy Murphy, Shane Usher and Jennifer Lane
Road(s) and routes per MDRS Map: South on Cow Dung Road, onto road 1101
Mode of travel: Curiosity Rover and Spirit Rover

Operations Report – November 19th

Crew 215 Operations Report 19-Nov-2019
SOL: 9
Name of person filing report: Shane Usher
Non-nominal systems: N/A
Notes on non-nominal systems: N/A
Generator: run
Hours run: 12 h
From what time last night: 1910
To what time this morning: 0710
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night): 78%
Diesel Reading – 46%
Station Propane Reading – 67%
Ethanol Free Gasoline: N/A
Water (loft tank) (gal): 42
Water Meter (units): 1459254
Water (static tank) (gal): 327
Static to Loft Pump used – yes
Water in Green Hab (gal): 243.7
Water in Science Dome (gal): 0
Toilet tank emptied: yes
Deimos rover used: no
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Sojourner rover used: Assigned to Director
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Spirit rover used: yes
Hours: 111.7
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 41%
Currently charging: yes
Opportunity rover used: no
Hours: 68
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes
Curiosity rover used: yes
Hours: 118.1
Beginning charge: (Before EVA) : 100%
Ending charge: (On return from EVA, before charging): 63%
Currently charging: yes
Notes on rovers: Spirit had high drain rate during use with 2 passengers.
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
ATV Fuel Used (gal): 0
# Hours the ATVs were used today: 0
Notes on ATVs: Assigned to the director
HabCar used and why, where? With Director
CrewCar used and why, where? At Grand Junction awaiting Crew 216
General notes and comments: N/A
Summary of internet: Functional
Summary of suits and radios: Voice Activation Transmission had been activated on some radios and caused general confusion. VOX now deactivated on all radios.
Summary of Hab operations: N/A
Summary of GreenHab operations: Atmospheric equalization to the green hab carried out during the afternoon, plants were watered (20L,5.7 gal). Door closed at 5pm, heater on for a short time.
Summary of Science Dome operations: Gypsum cement experiments conducted.
Summary of RAM operations: N/A
Summary of any observatory issues: Nominal
Summary of health and safety issues: All crew healthy
Questions, concerns and requests to Mission Support: Requested supplies arrived through RAM